WRITE (n-2)(3+(48/n))=48+22 in the standard form of a quadratic equation:
Answer by stanbon(75887)

WRITE (n-2)(3+(48/n))=48+22 in the standard form of a quadratic equation:
Let "n" be a common denominator in the 2nd factor of the 1st term.
------------------------
(n-2)[(3n+48)/n)] = 70
Multiply both sides by "n":
(n-2)(3n+48) = 70n
3n^2 + 48n - 6n - 96 = 70
3n^2 + 42n - 166 = 0
=============================
Cheers,
Stan H.